OpenStudy (narissa):

What is the slope of the line passing through the points (−1, 7) and (4, −1)? −2 −56 −85 2

OpenStudy (narissa):

@dumbcow

OpenStudy (eliesaab):

slope=\(\frac {y_2-y_1}{x_2 - x_1}\)

OpenStudy (eliesaab):

slope= \( \frac {-1-7}{4+1}=-\frac 8 5\)

OpenStudy (narissa):

thanks